### v1

# REN: Anatomically-Informed Mixture-of-Experts for Interstitial Lung Disease Diagnosis

Abstract
Mixture-of-Experts (MoE) architectures have significantly contributed to scalable machine learning by enabling specialized subnetworks to tackle complex tasks efficiently. However, traditional MoE systems lack domain-specific constraints essential for medical imaging, where anatomical structure and regional disease heterogeneity strongly influence pathological patterns. Here, we introduce Regional Expert Networks (REN), the first anatomically-informed MoE framework tailored specifically for medical image classification. REN leverages anatomical priors to train seven specialized experts, each dedicated to distinct lung lobes and bilateral lung combinations, enabling precise modeling of region-specific pathological variations. Multi-modal gating mechanisms dynamically integrate radiomics biomarkers and deep learning (DL) features (CNN, ViT, Mamba) to weight expert contributions optimally. Applied to interstitial lung disease (ILD) classification, REN achieves consistently superior performance: the radiomics-guided ensemble reached an average AUC of 0.8646 $\pm$ 0.0467, a +12.5% improvement over the SwinUNETR baseline (AUC 0.7685, $p=0.031$). Region-specific experts further revealed that lower-lobe models achieved AUCs of 0.88-0.90, surpassing DL counterparts (CNN: 0.76-0.79) and aligning with known disease progression patterns. Through rigorous patient-level cross-validation, REN demonstrates strong generalizability and clinical interpretability, presenting a scalable, anatomically-guided approach readily extensible to other structured medical imaging applications.
I Introduction
In machine learning, Mixture-of-experts (MoE) architectures include multiple expert networks that focus on different aspects of complex data distributions. The main principle underlying MoE success lies in conditional computation; dynamically routing inputs to the most relevant experts while maintaining computational efficiency. However, existing MoE frameworks primarily operate under unconstrained optimization structures that assume homogeneous expert capabilities and domain-agnostic routing mechanisms. We argue that this approach is fundamentally misaligned with medical imaging. Anatomical structure, physiological constraints, and regional disease heterogeneity impose domain-specific specialization requirements that generic routing cannot capture effectively.
Unlike natural image classification, where semantic regions may be arbitrarily distributed, medical images exhibit well-defined anatomical structure where pathological patterns follow predictable regional distributions governed by underlying physiological processes  [1]. Traditional MoE systems cannot leverage these domain-specific priors, leading to suboptimal expert utilization and reduced interpretability, a critical requirement for clinical adoption  [2].
For a specific example in automatic diagnosis of interstitial lung disease (ILD), one can easily exemplify these challenges. ILD encompasses over 200 diverse pulmonary disorders. Accurate diagnosis from high-resolution computed tomography (HRCT) scans is essential, since HRCT plays a central role in distinguishing ILD subtypes and guiding diagnosis in clinical practice [3]. Since each anatomical region may exhibit varying disease manifestations, severity levels, and progression patterns within the same patient, conventional global analysis approaches cannot effectively diagnose with interpretable computation. Current DL methods for ILD classification treat the entire lung as a homogeneous unit, potentially diluting region-specific pathological signals and providing limited interpretability for clinical decision-making [4].
Figure 1: Overview of the REN (Regional Expert Networks) framework. (A) Anatomical region extraction: preprocessing and lobe segmentation assign CT regions to seven experts (five lobes plus bilateral lungs). (B) Individual expert training: CNN, ViT, Mamba, and radiomics (XGBoost) experts are trained on masked inputs with validation AUCs recorded. (C) Gating function extraction: dynamic weighting strategies (performance-, feature-, and learned-based) are applied to expert outputs. (D) End-to-end MoE integration: expert weights and global SwinUNETR features are fused for patient-level ILD classification.
To this end, we introduce REN (Regional Expert Networks), a novel anatomically-informed MoE framework that addresses the aforementioned limitations by embedding domain-specific constraints into expert specialization and routing mechanisms (Fig. 1). Our novel approach demonstrates how anatomical priors can enhance MoE architectures. By training specialized experts for distinct anatomical regions and implementing multi-modal adaptive gating strategies, REN leverages inherent structural knowledge in medical applications while maintaining the scalability advantages of MoE systems. Through rigourous evaluation, we show that REN enables effective integration of both DL and traditional radiomics approaches. Particularly, radiomics-guided newly designed gating strategies prove effective for capturing subtle pathological patterns in anatomically critical regions. Our contributions can be summarized as follows:
  • Novel Anatomically-Constrained MoE Architecture: We introduce a new MoE framework specifically designed for medical imaging that incorporates anatomical structure as explicit constraints, addressing critical gaps in domain-specific expert routing. This is done by training seven lobe-specific experts that specialize in region-specific pathological pattern recognition through masked CT inputs.
  • Multi-Modal Gating Mechanisms: We develop new gating strategies that integrate both DL features and traditional radiomics biomarkers, to dynamically balance complementary information sources.
  • Comprehensive Architecture Evaluation: We systematically compare three state-of-the-art architectures—a custom 3D CNN [5], a 3D Vision Transformer adapted from Dosovitskiy et al. [6], and (Vision) Mamba [7] architecture—across all anatomical regions within the lung using rigorous patient-level cross-validation, establishing architecture-specific regional specialization insights.
  • End-to-End Trainable Framework: We implement a fully differentiable MoE system that enables joint optimization of expert networks and gating mechanisms while maintaining interpretability through expert contribution analysis.
II Related Work
II-A Mixture of Experts in Deep Learning
MoE architectures leverage routing, expert diversity, and load balancing to enable scalable, efficient learning, as detailed in recent surveys [8]. While foundational for complex tasks, current MoE systems lack the domain-specific constraints critical for medical imaging, where anatomical structure guides disease patterns. A core challenge remains balancing expert specialization with computational cost. DeepSeek [9] advanced scalable MoE training through fine-grained expert partitioning and shared experts for generalization, whereas MoE++ [10] introduced zero-computation experts to enable efficient dynamic routing.
Extensions to vision and multi-modal settings have made MoE increasingly relevant to healthcare. MoE-LLaVA [11] achieved sparse activation in vision-language models with only 3B active parameters, matching larger dense models via MoE-Tuning—a key step toward making MoE viable in resource-constrained clinical settings. Despite these advances, adapting MoE frameworks to incorporate anatomical priors and clinical semantics remains an open frontier for medical imaging.
Recent MoE research in medical imaging has also begun exploring domain-informed expert assignment and multi-modal integration. Wu et al. [2] proposed heterogeneous tissue experts for gigapixel whole-slide images and Jiang and Shen [12] introduced M4oE, a foundation model for multi-modal segmentation using MoE routing These efforts highlight growing interest in adapting MoE to biomedical applications.
Our work differs by introducing an anatomically-constrained MoE tailored specifically for systemic sclerosis-associated ILD. Unlike prior approaches that define experts by tissue type, modality, or distributional shifts, we explicitly align experts with anatomy regions—mirroring radiological practice and enabling interpretable, region-level contributions. Furthermore, we integrate handcrafted radiomics features with deep expert outputs, yielding hybrid gating strategies that couple domain priors with multi-modal signals. This anatomically- and clinically-grounded orientation distinguishes our study from broader MoE developments and addresses interpretability demands unique to medical imaging.
II-B Radiomics and Multi-Modal Integration
Radiomics remains a powerful and actively utilized methodology in medical imaging, particularly when integrated with DL techniques. By extracting quantitative features describing texture, shape, and intensity patterns, radiomics provides complementary information to the hierarchical representations learned by deep networks [13]. This synergy continues to demonstrate significant effectiveness in improving diagnostic accuracy through multi-modal fusion. Traditional machine learning classifiers like XGBoost [14] maintain relevance for high-dimensional radiomics tasks due to their interpretability and capacity to model complex feature interactions. Nevertheless, current fusion strategies often overlook regional anatomical heterogeneity and differential diagnostic importance across tissue regions. Anatomically-informed integration of radiomics with DL represents a promising yet underexplored frontier for enhancing both accuracy and clinical interpretability.
II-C Anatomical Specialization and Regional Analysis
Integrating anatomical priors into DL models enhances performance and interpretability by emulating clinical radiology’s regional examination practices [15]. Regional analysis proves particularly effective for pathologies with anatomically specialized patterns. Conventional approaches that rely on whole-organ analysis often assume anatomical homogeneity and risk overlooking localized variation in pathology [16]. In pulmonary imaging, for example, global feature extraction across entire lung volumes can obscure localized disease signals, reducing both sensitivity and interpretability [17]. This underscores the clinical need for anatomically informed models that capture region-specific patterns and align outputs with radiological workflows [18].
Recent advances incorporate anatomical knowledge through: (i) Denoising autoencoders and specialized architectures [19], (ii) Shape priors for segmentation accuracy [20], and (iii) Domain-informed constraints improving diagnostics [21]. Brain tumor segmentation exemplifies significant gains from anatomical priors [22]. This aligns naturally with MoE frameworks, where specialized subnetworks process distinct anatomical regions. The shift from generic to anatomically-informed expert routing represents a novel MoE application in medical imaging.
II-D Gaps and Motivation
Despite these advances, current MoE frameworks lack the domain-specific constraints essential for medical applications, where anatomical structure and regional heterogeneity fundamentally influence disease manifestation patterns. The combination of anatomical expertise with advanced gating mechanisms represents an underexplored opportunity to enhance both diagnostic accuracy and clinical interpretability in medical image analysis. Our work addresses these limitations by introducing the first MoE framework specifically designed for medical imaging that incorporates anatomical structure as explicit constraints, enabling effective integration of both DL and traditional radiomics approaches.
III Methods
Our proposed anatomically-informed mixture-of-experts framework for ILD diagnosis (classification) consists of a four-stage pipeline that progressively builds from individual expert training to gating function extraction and finally to end-to-end MoE model creation. Fig. 1 represents REN in four stages: (A) anatomical region extraction through preprocessing and segmentation, (B) training of regional experts using CNN, ViT, Mamba, and radiomics models, (C) derivation of gating strategies to dynamically weight experts, and (D) end-to-end MoE integration with SwinUNETR for ILD classification.
III-A Dataset
Our study utilized a retrospective dataset from the Northwestern Scleroderma Registry comprising 597 patients with 1,898 longitudinal chest CT scans acquired between 2001 and 2023. This study was approved by the Northwestern University Institutional Review Board. All participants provided informed consent for inclusion in the Northwestern Scleroderma Registry at the time of enrollment. The cohort included 489 (81.9%) female patients with a mean age of 63.7 ± 12.7 years (range: 22.1-98.3 years). The study population was predominantly White (479 patients, 80.2%), with 84 (14.1%) Black, 20 (3.4%) Asian, and 14 (2.3%) other race patients. The dataset focused on systemic sclerosis-related conditions with the following disease subtype distribution: limited cutaneous systemic sclerosis (lcSSc): 284 (47.6%), diffuse cutaneous systemic sclerosis (dcSSc): 245 (41.0%), systemic sclerosis sine scleroderma (SSS): 28 (4.7%), with other subtypes (6.7%). Of the total cohort, 365 patients (61.1%) had confirmed ILD, forming the positive class for our binary classification task. Patients contributed an average of 3.2 scans each, reflecting longitudinal monitoring patterns typical in systemic sclerosis care.
III-B Data Preprocessing and Evaluation Strategy
All CT scans, stored as NIFTI files, were processed using identical preprocessing pipelines, accompanied by corresponding lobe segmentation masks and binary classification labels. Lung and lobe segmentation were performed using the state-of-the-art lungmask algorithm [23], which provides robust automated segmentation of pulmonary lobes across diverse imaging protocols and pathological conditions. Our anatomical specialization approach focuses on seven distinct lung regions (automatically determined with lungmask segmentation: the five individual lobes (Left Upper Lobe - LUL, Left Lower Lobe - LLL, Right Upper Lobe - RUL, Right Middle Lobe - RML, Right Lower Lobe - RLL) plus two combined regions (Left Lung combining lobes 1 and 2, Right Lung combining lobes 3, 4, and 5) (Fig. 2). For evaluation, we implement strict patient-level cross-validation to prevent data leakage, where multiple scans per patient could artificially inflate performance if distributed across training and evaluation sets.
Each fold assigns 80% of patients to training and 20% to holdout, where the 20% holdout patients are unique to that fold. The holdout patients are then split 50/50 into validation (10% of total patients) and test (10% of total patients) sets (non-overlap). Across all 5 folds, every patient appears exactly once in the 20% holdout position, ensuring 50% of patients serve as validation data and 50% as test data across folds, with no patient contamination between sets.
Feature-space visualization
We qualitatively assessed regional separability by projecting radiomics feature vectors into two dimensions using t-SNE with standard settings and a fixed random seed. Embeddings were generated on the common set of scans across experts and are shown by anatomical region and by ILD status.
Figure 2: Anatomical region extraction pipeline showing the process of generating masked inputs for each of the seven lung regions. The original CT scan is processed with lobe-specific segmentation masks to create region-focused inputs that enable expert specialization.
III-C Regional Expert Networks (RENs) Architecture
Proposed REN modules specialize in distinct anatomical regions, each processing masked CT inputs corresponding to its assigned lung lobe or lung combination. This design ensures that learning is focused on region-specific pathological patterns rather than global averages. We implement several backbone variants: CNN RENs adopt a 3D convolutional architecture with three progressive convolutional blocks (Conv3D-1, Conv3D-2, Conv3D-3) followed by two fully connected layers (FC-1, FC-2) for binary classification. ViT RENs use a 3D Vision Transformer with patch embedding, positional encoding, multi-head transformer blocks, and a classification head. Mamba RENs apply a state-space model architecture with 3D patch embedding, stacked Mamba blocks with configurable state dimensions, global pooling, and linear classification layers. This multi-architecture design allows comparative evaluation of regional specialization across convolutional, attention-based, and state-space modeling structures.
III-D Radiomics RENs
For the five individual lung lobes and two lung regions, we extract comprehensive radiomics feature sets using PyRadiomics [24]. The radiomics pipeline computes first-order statistics (mean, variance, skewness, kurtosis), shape-based features (volume, surface area, sphericity, compactness), and texture features including Gray Level Co-occurrence Matrix (GLCM), Gray Level Run Length Matrix (GLRLM), Gray Level Size Zone Matrix (GLSZM), Gray Level Dependence Matrix (GLDM), and Neighbourhood Grey Tone Difference Matrix (NGTDM) features. This feature set captures complementary aspects of lung tissue characterization. First-order features quantify density and heterogeneity, shape features capture anatomical alterations, and texture features detect spatial and microstructural patterns associated with ILD-related pathophysiology. We used a 25 Hounsfield Unit bin width for intensity discretization to balance sensitivity to tissue differences while maintaining robustness to noise across different scanner protocols. This results in 107 quantitative features per lung lobe that capture subtle patterns potentially missed by DL approaches while providing clinically interpretable measurements. XGBoost classifiers are trained for each lobe and lung using these extracted features.
III-E Training Pipeline
Our end-to-end approach follows four sequential stages: (Stage 1) anatomical preprocessing and lobe segmentation; (Stage 2) regional expert training; (Stage 3) ensemble and gating extraction; and (Stage 4) end-to-end MoE optimization.
Stage 2 - Individual Regional Expert Network Training: In Stage 2, we trained specialized RENs for seven anatomical lung regions using three DL architectures (CNN, ViT, Mamba) and one radiomics model (XGBoost). Each REN received a masked CT input, isolating the target region via element-wise multiplication with its anatomical mask. Deep learning models were trained with architecture-specific hyperparameters and medical imaging–oriented augmentations (rotation, noise, affine transforms). CNN and Mamba used a learning rate of $1\times 10^{-4}$, batch size of 4; ViT used a learning rate of $5\times 10^{-5}$ and a batch size 2 due to memory demands. All models were optimized with AdamW and regularized with dropout, early stopping, and label smoothing. This stage yielded 28 RENs per fold (7 regions × 4 models), with validation AUCs used in Stage 2 ensemble integration.
Stage 3A - Basic Ensemble Evaluation: Stage 3 includes two sub-steps. Stage 3A establishes a simple ensemble baseline by combining the outputs of the regional experts rather than relying on any single model. The motivation for using an ensemble is that different experts capture complementary signals from distinct lung regions, so aggregating them can reduce variance and improve overall stability. To formalize this, we compute a weighted prediction for each input, where the contribution of each expert is scaled by its validation AUC. This produces a single ensemble output that reflects both the predictions and the relative reliability of all experts. For a given input sample $\mathbf{x}$ with corresponding lobe mask $\mathbf{m}$, each expert $E_{k}$ produces a prediction $\hat{y}_{k}(\mathbf{x})$, where $k\in\{1,2,\ldots,K\}$ represents the seven anatomical regions with $K=7$. Equation (1) defines the weighted ensemble prediction:
$\hat{y}_{\text{weighted}}(\mathbf{x})=\sum_{k=1}^{K}w_{k}\cdot\hat{y}_{k}(\mathbf{x}),$ (1)
where the normalized expert weights are computed as
$w_{k}=\frac{\text{AUC}_{k}^{\text{val}}}{\sum_{j=1}^{K}\text{AUC}_{j}^{\text{val}}}.$ (2)
Stage 3B - Advanced Gating Function Extraction: Stage 3B derives expert weighting strategies by analyzing validation behavior and internal representations of the trained RENs. Gating refers to the mechanism that decides how much influence each expert should have in the final ensemble. Unlike static averaging, gating adapts weights to reflect differences in expert reliability, patient-specific variation, or feature-level signal strength. This step is necessary because regional experts capture heterogeneous aspects of disease, and naïve equal weighting can obscure important patterns. By designing new gating strategies, we allow the ensemble to dynamically emphasize the most informative experts while downweighting those that contribute less useful or redundant information. We extract three gating function types: (1) Performance-based (6 total) using AUC normalization (3)–(5) and dynamic gating based on prediction confidence (6)–(8); (2) Feature-based (3 total), leveraging intermediate activations for input-adaptive weights via magnitude (9), variance (10), and entropy (11); and (3) Learned gating networks (2 total), training lightweight neural networks to predict expert weights with softmax (22) or sigmoid (23) normalization. Each REN type (CNN, ViT, Mamba) contributes identically structured gating functions to ensure fair comparison. For every model–fold pair (3 models × 5 folds), we evaluate all gating strategies and select the one yielding the best validation ensemble AUC. The result is a set of optimal gating weights $\mathbf{w}^{*}_{\text{model,fold}}\in\mathbb{R}^{7}$, used in Stage 3 to construct the final prediction ensemble.
III-E1 Enhanced AUC Normalization
Various normalization schemes, such as softmax, sigmoid, and sparsemax are applied to validation AUC scores in (3)–(5),
$g_{k}^{\text{val-auc-softmax}}$ $=\frac{\exp(\text{AUC}_{k}^{\text{val}})}{\sum_{j=1}^{K}\exp(\text{AUC}_{j}^{\text{val}})},$ (3)
$g_{k}^{\text{val-auc-sigmoid}}$ $=\frac{\sigma(\text{AUC}_{k}^{\text{val}})}{\sum_{j=1}^{K}\sigma(\text{AUC}_{j}^{\text{val}})},$ (4)
$g_{k}^{\text{val-auc-sparsemax}}$ $=\max(\text{AUC}_{k}^{\text{val}}-\tau,0),$ (5)
where $\sigma(x)=1/(1+e^{-x})$ and $\tau$ ensures unity sum for sparsemax. Softmax emphasizes small differences between AUC scores by producing dense probability distributions, sigmoid compresses the range of values to create smoother weightings, and sparsemax projects onto the simplex with exact zeros to encourage sparsity. These approaches differ in how strongly they amplify, smooth, or prune expert contributions.
III-E2 Dynamic Performance Gating
Sample-dependent weights based on prediction characteristics are defined in (6)–(8),
$g_{k}^{\text{confidence}}(\mathbf{x})$ $=\frac{\exp\!\left(2|\hat{y}_{k}(\mathbf{x})-0.5|\right)}{\sum_{j=1}^{K}\exp\!\left(2|\hat{y}_{j}(\mathbf{x})-0.5|\right)},$ (6)
$g_{k}^{\text{error}}(\mathbf{x})$ $=\frac{\exp\!\left(-|\hat{y}_{k}(\mathbf{x})-y|\right)}{\sum_{j=1}^{K}\exp\!\left(-|\hat{y}_{j}(\mathbf{x})-y|\right)},$ (7)
$g_{k}^{\text{diversity}}$ $=\frac{\exp(1-\bar{\rho}_{k})}{\sum_{j=1}^{K}\exp(1-\bar{\rho}_{j})},$ (8)
where $\bar{\rho}_{k}$ represents the average correlation between expert $k$ and all others. Confidence, error, and diversity each address complementary aspects of expert reliability. Confidence weighting gives more influence to experts that make decisive predictions far from uncertainty, error weighting prioritizes experts that consistently align with the ground truth, and diversity weighting favors experts that provide outputs less correlated with the others.
III-E3 Statistical Feature Analysis
Feature-vector-based weighting strategies are shown in (9)–(11),
$g_{k}^{\text{magnitude}}(\mathbf{x})$ $=\frac{\exp(\|\boldsymbol{\phi}_{k}(\mathbf{x})\|_{2})}{\sum_{j=1}^{K}\exp(\|\boldsymbol{\phi}_{j}(\mathbf{x})\|_{2})},$ (9)
$g_{k}^{\text{variance}}(\mathbf{x})$ $=\frac{\exp(\text{Var}(\boldsymbol{\phi}_{k}(\mathbf{x})))}{\sum_{j=1}^{K}\exp(\text{Var}(\boldsymbol{\phi}_{j}(\mathbf{x})))},$ (10)
$g_{k}^{\text{entropy}}(\mathbf{x})$ $=\frac{\exp(H_{k}(\mathbf{x}))}{\sum_{j=1}^{K}\exp(H_{j}(\mathbf{x}))},$ (11)
where $\boldsymbol{\phi}_{k}(\mathbf{x})$ represents extracted features and $H_{k}(\mathbf{x})=-\sum_{b=1}^{20}p_{k,b}\log(p_{k,b}+\epsilon)$ is entropy from 20-bin histograms. Magnitude weighting emphasizes experts with stronger overall feature activations, variance weighting favors experts whose features display greater spread, and entropy weighting highlights experts with more uniform distributions. These methods differ in whether they capture signal strength, variability, or uncertainty, providing complementary perspectives for weighting experts.
III-E4 Learned Gating Network Architectures
The CNN gating network applies multi-layer perceptron processing of FC1 features as in (12)–(14),
$\mathbf{h}_{1}$ $=\text{ReLU}(\mathbf{W}_{1}\overline{\boldsymbol{\phi}}^{\text{FC1}}(\mathbf{x})+\mathbf{b}_{1}),\quad\text{64 units},$ (12)
$\mathbf{h}_{2}$ $=\text{ReLU}(\mathbf{W}_{2}\mathbf{h}_{1}+\mathbf{b}_{2}),\quad\text{32 units},$ (13)
$\boldsymbol{z}(\mathbf{x})$ $=\mathbf{W}_{3}\mathbf{h}_{2}+\mathbf{b}_{3},\quad 7\text{ outputs}.$ (14)
The Mamba gating network incorporates multi-head attention in (15)–(17),
$\mathbf{h}_{\text{proj}}$ $=\mathbf{W}_{\text{proj}}\overline{\boldsymbol{\phi}}^{\text{pre-cls}}(\mathbf{x})+\mathbf{b}_{\text{proj}},$ (15)
$\mathbf{h}_{\text{attn}}$ $=\text{MultiHeadAttention}(\mathbf{h}_{\text{proj}},\mathbf{h}_{\text{proj}},\mathbf{h}_{\text{proj}}),$ (16)
$\boldsymbol{z}(\mathbf{x})$ $=\text{MLP}(\mathbf{h}_{\text{attn}}).$ (17)
The ViT gating network applies transformer-style processing with normalization as in (18)–(21),
$\mathbf{h}_{\text{proj}}$ $=\text{LayerNorm}(\mathbf{W}_{\text{proj}}\overline{\boldsymbol{\phi}}^{\text{CLS}}(\mathbf{x})+\mathbf{b}_{\text{proj}}),$ (18)
$\mathbf{h}_{\text{attn}}$ $=\text{LayerNorm}(\text{MultiHeadAttention}(\mathbf{h}_{\text{proj}})+\mathbf{h}_{\text{proj}}),$ (19)
$\mathbf{h}_{\text{mlp}}$ $=\text{LayerNorm}(\text{MLP}(\mathbf{h}_{\text{attn}})+\mathbf{h}_{\text{attn}}),$ (20)
$\boldsymbol{z}(\mathbf{x})$ $=\mathbf{W}_{\text{head}}\mathbf{h}_{\text{mlp}}+\mathbf{b}_{\text{head}}.$ (21)
The CNN gating network uses FC1 activations because they provide compact mid-level features, and a lightweight MLP can map them into expert weights efficiently. The Mamba gating network leverages state-space projection with multi-head attention to capture sequential structure and contextual dependencies in pre-classification features, which improves adaptive weighting. The ViT gating network applies transformer-style normalization and attention on the CLS token, enabling the gating module to model global context and assign expert weights in a more input-adaptive and context-aware manner.
Summary of Gating Network Formulations. Gating is the mechanism that determines how much each expert contributes to the final decision, and it is necessary because regional experts capture different aspects of disease and must be adaptively weighted rather than combined with fixed averages. The CNN gating network equations (12)–(14) specify a simple two-layer perceptron producing seven expert logits.
The Mamba gating network incorporates sequential feature processing through projection, multi-head attention, and an MLP as defined in (15)–(17). The ViT gating network follows a transformer-style design with layer normalization and residual connections, detailed in (18)–(21). Together, these formulations ensure that each architectural variant applies a consistent mapping from REN features to expert weight logits prior to normalization in (22) and (23).
Final gating weights for learned networks are normalized as in (22)–(23),
$g_{k}^{\text{learned-softmax}}(\mathbf{x})$ $=\frac{\exp(z_{k}(\mathbf{x}))}{\sum_{j=1}^{K}\exp(z_{j}(\mathbf{x}))},$ (22)
$g_{k}^{\text{learned-sigmoid}}(\mathbf{x})$ $=\frac{\sigma(z_{k}(\mathbf{x}))}{\sum_{j=1}^{K}\sigma(z_{j}(\mathbf{x}))}.$ (23)
III-E5 Gating Network Training Configuration
Learned gating networks were trained with AdamW (learning rate $0.001$, weight decay $10^{-4}$), early stopping (patience 15), and gradient clipping (max norm 1.0). The loss function is given in (24),
$\mathcal{L}_{\text{gating}}=\mathcal{L}_{\text{BCE}}(\hat{y}_{\text{ensemble}},y)+\lambda\mathcal{L}_{\text{entropy}}(\mathbf{g}),$ (24)
with $\lambda=0.01$.
Stage 4 - End-to-End MoE Architecture Training: The final stage trains a unified DL (specifically, we chose SwinUNETR as backbone) mixture-of-experts model for full lung classification, initialized with optimal gating strategies from Stage 3. Rather than training separate regional experts, Stage 4 uses a single end-to-end architecture that processes the entire lung while using anatomical expertise to weight regional contributions.
III-E6 Best Gating Strategy Selection
The best gating strategy for each expert type is selected according to Eq. (25),
$\mathbf{w}^{*}_{t}=\arg\max_{\mathbf{w}_{j}\in\mathcal{G}_{t}}\text{AUC}_{\text{val}}(\text{Ensemble}(\mathbf{w}_{j}),\mathcal{D}_{\text{val}}),$ (25)
where $\mathcal{G}_{t}$ represents the complete set of gating vectors for type $t$.
III-E7 Hierarchical Weight Normalization
Expert weights are normalized at lobe and lung levels as in (26)–(27),
$\mathbf{w}_{\text{lobes}}$ $=\frac{\mathbf{w}_{1:5}}{\sum_{i=1}^{5}w_{i}},$ (26)
$\mathbf{w}_{\text{lungs}}$ $=\frac{\mathbf{w}_{6:7}}{\sum_{i=6}^{7}w_{i}}.$ (27)
III-E8 Multi-Component Loss Function
To ensure that the end-to-end model learns not only accurate predictions but also balanced and interpretable expert contributions, we design a multi-component loss function that combines standard classification loss with additional regularizers for gating behavior, weight distribution, and expert diversity. The complete loss includes four components in (28)–(31),
$\mathcal{L}_{\text{CE}}$ $=-\tfrac{1}{N}\sum_{i=1}^{N}\!\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right],$ (28)
$\mathcal{L}_{\text{gating}}$ $=-\sum_{k=1}^{K}w_{k}\log(w_{k}+\epsilon),$ (29)
$\mathcal{L}_{\text{weight}}$ $=\|\mathbf{w}-\tfrac{1}{K}\mathbf{1}\|_{2}^{2},$ (30)
$\mathcal{L}_{\text{diversity}}$ $=\tfrac{1}{K^{2}}\sum_{i=1}^{K}\sum_{j\neq i}^{K}\!|\cos(\boldsymbol{\phi}_{i},\boldsymbol{\phi}_{j})|.$ (31)
Architecture Design: The proposed REN model processes the full CT volume $\mathbf{x}\in\mathbb{R}^{D\times H\times W}$ and lobe mask $\mathbf{m}\in\mathbb{R}^{D\times H\times W}$ using three main components. The SwinUNETR backbone extracts global features $\mathbf{F}_{\text{backbone}}=\text{SwinUNETR}(\mathbf{x})$, globally pooled to $\boldsymbol{\phi}_{\text{global}}\in\mathbb{R}^{384}$. Seven expert extractors process masked inputs $\mathbf{x}_{k}=\mathbf{x}\odot\mathbf{m}_{k}$ to produce features $\boldsymbol{\phi}_{k}\in\mathbb{R}^{64}$, with attention weights $\alpha_{k}$. Final expert weights are $g_{k}=w_{k}\alpha_{k}$, normalized such that $\sum_{k=1}^{K}g_{k}=1$. Weighted expert features $\boldsymbol{\phi}_{\text{expert}}=\sum_{k=1}^{K}g_{k}\boldsymbol{\phi}_{k}$ are concatenated with global features to form $\boldsymbol{\phi}_{\text{combined}}=[\boldsymbol{\phi}_{\text{global}};\boldsymbol{\phi}_{\text{expert}}]\in\mathbb{R}^{448}$, classified by an MLP as $y_{\text{pred}}=\text{MLP}(\boldsymbol{\phi}_{\text{combined}})$.
Weight Initialization Strategies: Stage 4 initializes and compares seven strategies: four based on validation AUC weighting from Stage 2 (radiomics, CNN, ViT, Mamba) and three based on the optimal learned gating from Stage 3 ($\mathbf{w}^{*}_{\text{CNN}},\mathbf{w}^{*}_{\text{ViT}},\mathbf{w}^{*}_{\text{Mamba}}$). Optimal weights are selected via (25), and expert weights are hierarchically normalized per Eqs. (26)–(27).
Training Configuration: The end-to-end SwinUNETR MoE is trained with AdamW using differentiated learning rates ($10^{-5}$ for backbone, $10^{-4}$ for MoE) and cosine annealing scheduling. The total loss is
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda_{\text{gating}}\mathcal{L}_{\text{gating}}+\lambda_{\text{weight}}\mathcal{L}_{\text{weight}}+\lambda_{\text{diversity}}\mathcal{L}_{\text{diversity}},$ (32)
with hyperparameters $\lambda_{\text{gating}}=0.005$, $\lambda_{\text{weight}}=0.005$, and $\lambda_{\text{diversity}}=0.01$ selected via validation.
III-F Summary of Methodological Innovations
First, we introduce RENs, the first anatomically informed MoE architecture in medical imaging, specifically in lung diseases. Expert assignment is guided explicitly by lobe- and lung-level segmentation rather than unconstrained routing. This ensures that expert specialization mirrors radiological practice and directly encodes anatomical priors into the model.
Second, we propose a novel multi-modal gating framework that integrates three complementary strategies: (i) performance-based weighting derived from validation AUCs, (ii) feature-based weighting using intermediate representations, and (iii) lightweight learned gating networks. This design enables dynamic, input-adaptive expert selection while avoiding the interpretability and load-balancing limitations of conventional MoE routing.
Third, we incorporate radiomics-guided expert models, providing handcrafted, pathology-aware biomarkers that complement deep learning features. By embedding radiomics in both regional experts and gating mechanisms, our framework enhances interpretability and sensitivity to subtle, anatomically localized disease patterns.
Collectively, these innovations constitute a domain-specific adaptation of MoE tailored for ILD classification. REN is not a simple application of existing MoE systems; it is a redesign that leverages anatomical priors, radiomics integration, and hybrid gating to produce clinically interpretable predictions.
IV Experiments and Results
Our anatomically-informed mixture-of-experts framework for ILD classification was evaluated through comprehensive cross-validation experiments across three architectures and multiple ensemble strategies. Results demonstrate the effectiveness of regional specialization and gating mechanisms for improving diagnostic accuracy while providing enhanced interpretability.
Model-Specific Hyperparameters: All DL experts were trained using model-specific hyperparameters optimized for medical imaging tasks. CNN and Mamba experts used learning rate $1\times 10^{-4}$, batch size 4, while ViT experts used learning rate $5\times 10^{-5}$, batch size 2. All models used AdamW optimizer with weight decay $1\times 10^{-5}$. Early stopping was applied with a patience of 15 epochs. Learning rate scheduling used ReduceLROnPlateau for CNNs and CosineAnnealingLR for Mamba and ViT models. Dropout rates were 0.5 for CNNs, 0.2 for Mamba and ViT. Label smoothing was 0.05 for CNN, 0.2 for ViT and Mamba.
Computational Environment and Reproducibility: All experiments were conducted on NVIDIA A100-SXM4-80GB GPUs, using CUDA 12.1 and driver version 525.85.12. The software environment consisted of Python 3.8.20, PyTorch 2.4.1+cu121, MONAI 1.3.2, NumPy 1.24.4, scikit-learn 1.3.2, Pandas 2.0.3, XGBoost 2.1.4, and pyradiomics v3.1.0. Mamba models used mamba-ssm 2.2.2 for state-space computations.
For reproducibility, all random number generators were initialized with seed=42 using a seeding function that sets Python’s random module, NumPy, PyTorch CPU and GPU generators, and enables CUDA deterministic operations. Cross-validation splits used fold-specific seeds to ensure consistent patient partitioning across all experiments.
IV-A Stage 2: Individual Expert Performance
Figure 3: Mean AUC per lung region across folds for each architecture. Radiomics experts achieved the highest regional performance, particularly in the lower lobes, while CNN followed closely behind. Mamba and ViT showed more variability across regions.
Architecture Performance Comparison: Radiomics-based MoE models outperformed neural networks based MoEs in most region-fold comparisons: 23/25 vs. Mamba, 24/25 vs. ViT, and 17/25 vs. CNN, with mean AUC differences of 0.117 ± 0.044 (95% CI: [0.073, 0.162]), 0.193 ± 0.053 (95% CI: [0.140, 0.246]), and 0.069 ± 0.046 (95% CI: [0.022, 0.115]), respectively (Fig. 3). This performance advantage likely stems from radiomics’ domain-specific features (n=107), which encode texture, shape, and intensity, offering compact, pathology-aware representations and reducing overfitting risks on limited data. Among neural networks, CNNs showed the closest performance to radiomics, suggesting convolutional filters effectively model local lung structures. More advanced models, ViT and Mamba, originally optimized for non-medical domains, lagged behind—highlighting the need for adaptation when applying transformer-style models to clinical imaging.
Regional Specialization Patterns: Model performance mirrored known ILD progression patterns, with highest AUCs in the left and right lower lobes (LLL/RLL), consistently exceeding 0.85 in the radiomics-based MoE approach. This aligns with clinical evidence that ILD typically begins in the lung bases due to gravitational and mechanical stress. Radiomics was particularly sensitive to early fibrotic features—honeycombing, reticulation, and ground-glass opacities—in these regions. In contrast, reduced performance in the upper lobes and right middle lobe (RML) across all methods reflects the later onset of ILD in these areas. These results suggest the models capture the true anatomical progression of ILD, rather than spurious spatial patterns.
IV-B Stage 3: Gating Strategy Evaluation
Learned vs. Static Gating Performance: Learned gating consistently outperforms static weighting across architectures, improving model stability through adaptive expert selection. This validates dynamic gating as essential for medical MoE architectures, where input-dependent routing captures patient-specific anatomical variations while maintaining predictable behavior across diverse cases.
IV-C Stage 4: End-to-End Integration Results
Tables I and II compare final test performance between weighted ensemble approaches, end-to-end MoE architectures, and baseline methods using different anatomical configurations.
TABLE I: Final Performance Using Five Individual Lung Lobes: Baseline vs. Validation-Weighted vs. End-to-End Gated MoE
Method AUC $\pm$ SD [95% CI] Change p-value
Baseline
SwinUNETR (Baseline) 0.7685 $\pm$ 0.0759 [0.674, 0.863]
Validation-Weighted
CNN Weighted 0.8033 $\pm$ 0.0472 [0.745, 0.862] +4.5% 0.176
ViT Weighted 0.8015 $\pm$ 0.0453 [0.745, 0.858] +4.3% 0.179
Mamba Weighted 0.7995 $\pm$ 0.0466 [0.742, 0.857] +4.0% 0.213
Radiomics-Guided 0.8646 $\pm$ 0.0467 [0.806, 0.923] +12.5% 0.031
End-to-End Gated MoE
CNN Gated MoE 0.7760 $\pm$ 0.0689 [0.691, 0.862] +1.0% 0.819
ViT Gated MoE 0.8010 $\pm$ 0.0430 [0.748, 0.854] +4.2% 0.166
Mamba Gated MoE 0.8164 $\pm$ 0.0393 [0.768, 0.865] +6.2% 0.053
TABLE II: Final Performance Comparison Using 5 Lobes + 2 Bilateral Lung Regions (7 Total): Baseline vs. Weighted MoE vs. Gated MoE
Method AUC $\pm$ SD [95% CI] Change p-value
Baseline
SwinUNETR (Baseline) 0.7685 $\pm$ 0.0759 [0.674, 0.863]
Validation-Weighted
CNN Weighted 0.7960 $\pm$ 0.0512 [0.732, 0.860] +3.6% 0.382
ViT Weighted 0.8021 $\pm$ 0.0577 [0.730, 0.874] +4.4% 0.297
Mamba Weighted 0.7983 $\pm$ 0.0490 [0.737, 0.860] +3.9% 0.351
Radiomics-Guided 0.8523 $\pm$ 0.0430 [0.796, 0.908] +10.9% 0.014
End-to-End Gated MoE
CNN Gated MoE 0.7778 $\pm$ 0.0545 [0.710, 0.846] +1.2% 0.442
ViT Gated MoE 0.7946 $\pm$ 0.0328 [0.754, 0.835] +3.4% 0.242
Mamba Gated MoE 0.7950 $\pm$ 0.0486 [0.735, 0.855] +3.4% 0.278
Anatomical Configuration Comparison: We evaluated two anatomical expert configurations: (1) five individual lung lobes only, and (2) seven regions including five individual lobes plus bilateral lung combinations. The five-lobe configuration achieved superior performance, with the radiomics-guided ensemble reaching an average AUC of 0.8646 compared to 0.8523 for the seven-region configuration. This 1.4% improvement suggests that including bilateral lung experts may introduce redundancy that dilutes the specialized knowledge of individual lobe experts.
Primary Performance and Computational Burden: The radiomics-guided ensemble achieved the highest performance using the five-lobe configuration (average AUC 0.8646). This represents a statistically significant 12.5% improvement over baseline SwinUNETR ($p=0.031$, paired t-test across CV folds). This demonstrates the complementary value of engineered features when combined with anatomically informed expert weighting. All MoE variants meet typical clinical deployment requirements. Inference times are 46–48 ms per patient (21+ samples/s) for DL approaches and 62 ms for radiomics-guided ensembles. Model sizes (237–247 MB; 62–65 M parameters) remain manageable for standard clinical hardware. These speeds far exceed human interpretation time, enabling real-time diagnostic support. Importantly, weighted ensembles consistently outperform end-to-end gated counterparts. This validates our training strategy. This staged approach—training experts independently before combination—preserves specialized anatomical knowledge that joint optimization obscures. The modest computational overhead of radiomics integration ($\approx 16$ ms) is negligible relative to its 12.5% performance gain. This establishes the hybrid approach as both computationally feasible and diagnostically superior for clinical deployment.
Clinical Translation and Significance: The 12.5% improvement could translate to better identification of early-stage ILD cases, where timely diagnosis is critical for patient outcomes. The statistical significance of only the radiomics-guided approach suggests that while other methods show numerical improvements, they may not provide reliable enough enhancement for clinical implementation.
Figure 4: Radiomics feature category ablation analysis showing (left) performance using individual feature categories across all lung regions and (right) feature importance measured by AUC drop when each category is removed. Texture features demonstrate particular strength in lower lobe regions, with GLCM achieving 0.828 AUC in Left Lower Lobe and GLRLM reaching 0.845 in Right Lower Lobe.
Radiomics Analysis: Comprehensive ablation studies reveal texture features (GLCM, GLRLM) drive performance in lower lobes where ILD initiates (Fig. 4). The embeddings corroborate the quantitative results: texture families (GLCM/GLRLM) yield tight, separable lower-lobe clusters, while upper-lobe geometry remains fragmented, explaining regional AUC differences. Lower lobe clustering patterns align with clinical knowledge of disease progression, while the approach’s consistency across cross-validation folds ensures reliability across diverse clinical scenarios—critical for real-world deployment. The 7.6-8.1% improvement over individual ensembles demonstrates tangible workflow benefits. This hybrid approach combines radiomics interpretability with DL’s pattern recognition, providing both diagnostic accuracy and explainable insights. Regional performance breakdowns enable targeted radiologist attention and support differential diagnosis based on anatomical distribution patterns, aligning with natural radiological evaluation practices and facilitating clinical integration.
V Discussion and Concluding Remarks
In conclusion, REN establishes a novel paradigm for anatomically-informed MoE architectures in medical imaging, demonstrating superior ILD classification through lobe-specific expert specialization and multi-modal gating that effectively integrates radiomics and deep learning features. By achieving AUCs of 0.88–0.90 in lower lobes—outperforming clinical benchmarks and traditional DL approaches—REN addresses critical gaps in regional disease heterogeneity, providing interpretable insights via gating weights that align with clinical patterns of basal fibrosis. This framework’s scalability, validated through rigorous patient-level cross-validation, positions it as a transferable solution for other anatomically complex tasks, such as multi-subtype ILD diagnosis or multi-organ segmentation.
Domain-Informed MoE and Methodological Innovations: Our REN introduces a domain-specific Mixture-of-Experts framework that explicitly embeds anatomical priors into the architecture—a departure from traditional, unconstrained MoE designs. By aligning expert assignment with known lung regions, REN ensures that anatomical structure guides learning from the outset, rather than emerging through post-hoc interpretation. This yields interpretable regional contribution weights, enabling clinicians to validate model reasoning and relate predictions to established radiological workflows.
We contribute three core methodological innovations. First, cross-architecture analysis of CNN, ViT, and Mamba experts reveals region-specific strengths, suggesting that heterogeneous expert assignment may outperform uniform deployment in anatomically diverse tasks. Second, our evaluation of 11 gating strategies establishes learned gating as consistently superior (1.0–6.2% AUC gain), particularly for ViT, where attention-based mechanisms mitigate domain misalignment. Third, integration of radiomics—handcrafted features rooted in clinical imaging biomarkers—yields a 12.5% performance boost, especially in lower lobes where ILD first manifests. This synergy highlights how radiomics captures subtle, structured signals often missed by end-to-end networks.
Collectively, REN bridges the gap between DL and clinical intuition, transforming black-box predictions into anatomically grounded insights. By coupling domain priors with multi-modal fusion and dynamic gating, our framework advances medical image analysis toward more interpretable, effective, and clinically actionable AI systems.
Radiomics-Guided Hybrid Learning: Our four-stage pipeline demonstrates radiomics’ advantages for anatomically-informed modeling. Lower lobes consistently outperformed upper lobes by 0.068–0.114 AUC, mirroring ILD pathophysiology, with texture features (GLCM, GLRLM) driving performance where disease initiates. Feature visualization shows compact ILD-positive clusters in lower lobes, validating radiomics’ ability to capture subvisual changes. The compact 107-feature space minimizes overfitting while preserving interpretability, with radiomics-enhanced ensembles improving AUC by 7.6–8.1% over pure DL models through design that aligns with radiological practice.
Limitations and Future Directions: Our study, while demonstrating strong performance, has several limitations. First, the dataset originates from a single institution and focuses exclusively on systemic sclerosis-related ILD, which may limit generalizability to other ILD subtypes and populations. The 2001–2023 temporal range also introduces variability in imaging protocols and scanner hardware. Additionally, reliance on automated lungmask segmentation may propagate boundary errors throughout the pipeline. From a methodological perspective, our current binary ILD classification does not capture the heterogeneity across ILD subtypes. Expert assignment is anatomically fixed, potentially limiting adaptability to disease-specific spatial patterns. The ensemble’s computational cost—due to training and deploying seven regional networks—may hinder scalability in clinical settings.
Because our framework is tailored to systemic sclerosis-related ILD, a rare autoimmune disease, and relies on anatomically-defined expert assignment, identifying suitable external datasets was a challenge. Publicly available ILD cohorts lack the necessary regional annotations, or focus on different disease etiologies, which constrains opportunities for straightforward replication. This strengthens internal validity but also highlights the need for carefully curated, multi-institutional collaborations to assess whether the observed performance gains translate across patient populations and imaging environments.
To address these challenges, we outline several future directions: (1) multi-institutional validation to assess robustness across demographics and imaging variations; (2) expansion to multi-class ILD classification; (3) efficient knowledge distillation strategies to compress the ensemble while retaining the observed 12.5% performance gain; and (4) adaptive expert assignment driven by pathological patterns rather than fixed anatomical regions. Incorporating graph-based MoE models, such as Graph Mixture of Experts (GMoE) [25], could enable relational routing among lung regions and capture anatomical adjacency or disease spread dynamics. In parallel, recent advances in modality-specialized MoE architectures such as MedMoE [26] highlights opportunities for extending anatomically informed frameworks into multimodal and vision-language domains, where REN could integrate imaging with radiology reports or other EHR modalities. These steps will support transition from proof-of-concept to clinically deployable AI systems, optimized for real-world variability and constrained environments.